OpenStudy (anonymous):

Consider the line y= 8x-4 . Find the equation of the line that is parallel to this line and passes through the point (3,-6) . Find the equation of the line that is perpendicular to this line and passes through the point (3,-6).

OpenStudy (anonymous):

you know that the equation of parallel lines are similar. the only difference is the y-intercept so the equation of the required parallel line will be y = 8x + c now as the line passes through (3, -6) it will satisfy the equation so -6 = 8(3)+c c = -6-24 c= -30 so the required equation is y = 8x -30

OpenStudy (anonymous):

now for perpendicular lines, the slope of the required line multiplied by the slope of the given line = -1 from here find the slope of the required line you have slope and one point, so you can find the equation of the line